Which linear function represents the line given by the point-slope equation y – 8 =1/2 (x – 4)?
Answer by fractalier(6550)

y – 8 = 1/2(x – 4)
y - 8 = (1/2)x - 2
y = (1/2)x + 6
You didn't put the choices here, so I am guessing it is this...